OpenStudy (help!!!!):

Help me graph the line through the point P having slope m. Find the coordinates of two other points on the line. 34. P(-3,-1), m=0.75

OpenStudy (help!!!!):

@Loser66 @SolomonZelman @zepdrix @Hero @Jaynator495

zepdrix (zepdrix):

Since we're given a `slope` and a `point`, it's probably best to write the equation for our line in `point slope form`.

zepdrix (zepdrix):

So here is what point-slope looks like:\[\large\rm y-\color{orangered}{y_o}=m(x-\color{orangered}{x_o})\]Where \(\large\rm (\color{orangered}{x_o},~\color{orangered}{y_o})\) is the given point.

zepdrix (zepdrix):

Plugging in our values gives us,\[\large\rm y-\color{orangered}{(-3)}=0.75(x-\color{orangered}{(-1)})\]Simplifies to,\[\large\rm y+3=0.75(x+1)\]

zepdrix (zepdrix):

This will help us find other points on the line.

OpenStudy (help!!!!):

nice! Thanks for the clarification!

zepdrix (zepdrix):

To graph it easily, we just need one more point. Then we can draw a line through the two points.

OpenStudy (help!!!!):

Yup I figured. I just got confused when it asked me to find another sets of points

zepdrix (zepdrix):

Ahh woops! I plugged the numbers in backwards!! :O Notice I plugged the -3 into the y.

zepdrix (zepdrix):

\[\large\rm y+1=0.75(x+3)\]My bad >.< Our point was this: \(\large\rm P(-3,-1)\) It was NOT this \(\large\rm (-1,-3)\) which is what I plugged in at the start.

zepdrix (zepdrix):

So we need another point? Hmm... Let's plug in x=0 and see if we can come up with something.

zepdrix (zepdrix):

\[\large\rm y+1=0.75(0+3)\]So we're currently working on finding this point: \(\large\rm (0,?)\) We're choosing 0 for x, and we're going to solve this equation for y to see the corresponding coordinate value.

zepdrix (zepdrix):

Understand how to solve for y in that setup? :o

OpenStudy (help!!!!):

Yuup! Also could you help me out with a word problem related to slope?

OpenStudy (help!!!!):

A ramp to provide handicapped people access to a certain building is to be constructed with a slope of 5%. If the entrance to the building is 3 ft above ground level, how long should the base of the ramp be?

zepdrix (zepdrix):

\(\large\rm m=5\text{%}\) So it would be nice if we could write this as an actual number, not a percentage. Do you remember how to convert it? :)

OpenStudy (help!!!!):

hmm is it out of 100? I have no clue

zepdrix (zepdrix):

Yes, `per` `cent` means `per` `100`. Think of the word `century`, which measures 100 years. So 5% means, 5 per 100. So we can write that as a fraction: \(\dfrac{5}{100}\)

OpenStudy (help!!!!):

so thats the slope?

zepdrix (zepdrix):

\[\large\rm m=\frac{5}{100}\]Yes, that's the slope. Let's simplify it down. Both top and bottom are divisible by 5.

OpenStudy (help!!!!):

so 1/20

zepdrix (zepdrix):